OpenStudy (anonymous):

in the second quadrant of a unit circle, sin=1/3, what is sec=?

OpenStudy (stamp):

sin = opposite side / hypotenuse sin = 1 /3 opposite side / hypotenuse = 1 / 3 opposite side = 1 hypotenuse = 3 

OpenStudy (stamp):

adjacent side = x opposite side = y hypotenuse = h x^2 + y^2 = h^2 x^2 + (1)^2 = (3)^2 x^2 + 1 = 9 x^2 = 8 x = sqrt(8) sec = hypotenuse / adjacent sec = 3 / sqrt(8) sec = 3sqrt(8) / 8 sec = 3 2sqrt(2) / 8 sec = 3sqrt(2)/4

OpenStudy (stamp):

Since this is in the second quadrant, your adjacent side should be flipped. So adjacent will have a negative value, and sec = -3sqrt(2)/4

OpenStudy (anonymous):

THANKS! and the sec would be?

OpenStudy (anonymous):

*tan

OpenStudy (stamp):

tan = opposite / adjacent

OpenStudy (anonymous):

Thanks!!!!!